# Extremal metrics involving scalar curvature

### v3

@@ 1 Introduction
For the case of $M=M_{0}$ and $f=\mathrm{id}_{M}$, he also gave the following type of rigidity theorem.
@@
As noted in [10], one can apply the construction of Lohkamp [23] to see that not all metrics on $M$ are area-extremal (for this definition, see the beginning of Section 2) if $\mathrm{dim}M\geq 3$. However, such an example is not given in an explicit way. That is, we can deduce the existence of such a metric but we cannot know any concrete properties involving its curvatures in general. In light of the above, this study aims to investigate relations between some rigidity phenomena involving scalar curvature and standard metrics in various senses. In particular, in this paper, we give some necessary conditions for metrics to be extremal in some senses (see Definitions 1.1, 1.2 and 1.3 below and the corollaries following them).
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying
$\begin{split}&~~~\frac{3n-2}{2n}\Delta_{g}R_{g}+\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{2}}{n}>R_{g}\cdot\max_{v\in T_{x}M,\,|v|_{g}=1}\mathrm{Ric}_{g}(v,v)~~\mathrm{on}~M,\\ &\left(\mathrm{resp.}\,\frac{3n-2}{2n}\Delta_{g}R_{g}+\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{2}}{n}>R_{g}\cdot\min_{v\in T_{x}M,\,|v|_{g}=1}\mathrm{Ric}_{g}(v,v)~~\mathrm{on}~M\right)\end{split}$ (1)
where $\overset{\circ}{\mathrm{Ric}_{g}}:=\mathrm{Ric}_{g}-\frac{R_{g}}{n}g$ is the traceless Ricci tensor. Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
$R_{g_{s}}>R_{g}\cdot\|g\|^{2}_{1,g_{s}}~~(\mathrm{resp.}~R_{g_{s}}<R_{g}\cdot\|g\|^{2}_{1,g_{s}})$ (2)
at each point of $M$. Here, $\|g\|_{1,g_{s}}:M\rightarrow[0,\infty)$ is the function on $M$ defined by
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying the following. Let $\lambda_{Ric,1}(x)\leq\cdots\leq\lambda_{Ric,n-1}(x)\leq\lambda_{Ric,n}(x)$ be the ordered eigenvalues of $\mathrm{Ric}_{g}$ on $T_{x}M~(x\in M)$.
  • $R_{g}\leq 0$ on $M$ and on $M$,
    $\begin{split}&~~~~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,1}>0\\ &\left(\mathrm{resp.}~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n}>0\right)\end{split}$
    or
  • $R_{g}\geq 0$ on $M$ and on $M$,
    $\begin{split}&~~~~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,n}>0.\\ &\left(\mathrm{resp.}~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{3n-2}{2n}\Delta_{g}R_{g}+\frac{R_{g}^{2}}{n}-R_{g}\cdot\lambda_{Ric,1}>0\right)\end{split}$
Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying that
$\left(\frac{3n-2}{2n}\Delta_{g}R_{g}+\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{2}}{n}\right)\cdot g>R_{g}\cdot\mathrm{Ric}_{g}~\mathrm{on}~TM.$
Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
$R_{g_{s}}\cdot g_{s}>R_{g}\cdot g~~(\mathrm{resp.}~R_{g_{s}}\cdot g_{s}<R_{g}\cdot g)$ (7)
at each point of $M$.
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ satisfying that
$\left(\frac{3n-2}{n}R_{g}\,\Delta_{g}R_{g}+2R_{g}\,\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}+\frac{R_{g}^{3}}{n}\right)\cdot g>R^{2}_{g}\cdot\mathrm{Ric}_{g}~\mathrm{on}~\Lambda^{2}TM.$
Then, there is a small constant $s>0$ (resp. $s<0$) depending only on $n,M$ and $g$ such that $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$ is a Riemmanian metric on $M$ and that
$R^{2}_{g_{s}}\cdot g_{s}>R^{2}_{g}\cdot g~~(\mathrm{resp.}~R^{2}_{g_{s}}\cdot g_{s}<R^{2}_{g}\cdot g)$ (8)
@@
  • All the assumptions of the above Theorems 1.3, 1.4, 1.5 and 1.6 especially imply that $R_{g}$ is not a constant on $M$ or
@@
  • If $\Delta_{g}R_{g}=0$ on $M$, then $R_{g}$ is sign-changing otherwise $R_{g}\equiv\mathrm{const}$ on $M$.
  • Interestingly, Dahl–Kröncke [8, 17] recently also discovered a relation between stability of Einstein metrics and certain type of scalar curvature rigidity. Since every Einstein metric does not satisfy any of the above assumptions (see Remark 1.2 above), our theorems above cannot be applied to Einstein metrics.
@@
By taking the contrapositive of each of the above theorems, we obtain necessary conditions for metrics to be scalar curvature rigid in the sense of the above definitions.
Corollary 1.1.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type I scalar curvature rigid in the sense of Listing, then
$\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}\leq R_{g}(x)\cdot\max_{v\in T_{x}M,\,|v|_{g}=1}\mathrm{Ric}_{g}(v,v)$
at some point $x\in M$.
Corollary 1.2.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type II scalar curvature rigid in the sense of Listing, then either of the following holds:
  • $\max_{M}R_{g}>0$ or
    $\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}-R_{g}(x)\cdot\lambda_{Ric,1}(x)\leq 0$
    at some point $x\in M$, and
  • $\min_{M}R_{g}<0$ or
    $\frac{3n-2}{2n}\Delta_{g}R_{g}(x)+\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}(x)+\frac{R_{g}^{2}(x)}{n}-R_{g}(x)\cdot\lambda_{Ric,n}(x)\leq 0$
    at some point $x\in M$.
Corollary 1.3.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type III scalar curvature rigid in the sense of Listing, then
$\left(\frac{3n-2}{n}\Delta_{g}R_{g}(x)+2\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}(x)+\frac{R_{g}^{2}(x)}{n}\right)\cdot g_{x}(v,w)\leq R_{g}(x)\cdot\mathrm{Ric}_{g}(x)(v,w)$
at some point $x\in M$ and some vectors $v,w\in T_{x}M$.
Corollary 1.4.
Let $M^{n}$ be a closed manifold of dimension $n\geq 2$. If a metric $g$ on $M$ is type IV scalar curvature rigid in the sense of Listing, then
$\begin{split}&\left(\frac{3n-2}{n}R_{g}(x)\,\Delta_{g}R_{g}(x)+2R_{g}(x)\,\|\overset{\circ}{\mathrm{Ric}_{g}}\|^{2}_{g}(x)+\frac{R_{g}^{3}(x)}{n}\right)\cdot g_{x}(v,w)\\ &\leq R^{2}_{g}(x)\cdot\mathrm{Ric}_{g}(x)(v,w)\end{split}$
at some point $x\in M$ and some two-vectors $v,w\in\Lambda^{2}T_{x}M$.
This paper is organized as follows. In Section 2, we review some rigidity results for certain Riemannian metrics with positive scalar curvature and a relation between “scalar minimum functional” and an extremal metric (see Remark 2.1). In Section 3, we describe a formula that is necessary to prove our main theorems. Furthermore, we consider statements of the same type as our main theorems on compact manifolds with boundary and non-compact complete manifolds. In Section 4, we prove our main theorems. In Section 5, we give some examples that satisfy the assumptions of Theorem 1.3, 1.4, 1.5 and 1.6. In Section 6, we present some further questions related to our main theorems. In Section 7, we give a proof of the formula in Section 3.
@@ 4 Proofs of Main Theorems
is orthogonal to the subspace $C^{\infty}(M)\cdot g$ (with respect to $L^{2}(M,g)$ inner product). Moreover, if $R_{g}$ is constant, then it is also a tt-tensor. Indeed, from the contracted second Bianchi identity:
$\delta\mathrm{Ric}_{g}=-\frac{1}{2}\nabla R_{g},$
and $R_{g}\equiv\mathrm{const}$, $\overset{\circ}{\mathrm{Ric}_{g}}$ is divergence-free. Since $\overset{\circ}{\mathrm{Ric}_{g}}$ is also trace-free, hence it is a tt-tensor. We are going to take this tensor $\overset{\circ}{\mathrm{Ric}_{g}}$ as the variation $h$ in the following proof of our main theorems.
@@
From the formula (13), we have for $0<s<<1$,
@@
$\begin{split}&R_{g-s(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g)}-R_{g}\cdot\|g\|^{2}_{1,\,g-s(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g)}\\ &=R_{g}+\frac{n-1}{n}s\Delta_{g}R_{g}-\frac{R_{g}^{2}}{n}s\\ &~~~~+\frac{s}{2}\Delta_{g}R_{g}+s\|\mathrm{Ric}_{g}\|^{2}_{g}-\frac{R_{g}}{1-s\overline{\lambda}_{Ric}+\frac{R_{g}}{n}s}\\ &+g_{s}^{-1}\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}*\mathrm{Ric}_{g}\\ &+Q_{s}(\nabla^{g}\mathrm{Ric}_{g}),\end{split}$
where $g_{s}:=g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)$. We have used here that
  • $DR|_{g}(-\mathrm{Ric}_{g})=\frac{1}{2}\Delta_{g}R_{g}+\|\mathrm{Ric}_{g}\|_{g}^{2}$, and
  • $DR|_{g}\left(R_{g}\right)=-(n-1)(\Delta_{g}R_{g})g$.
The first one is known as the first variation of the scalar curvature functional along the Ricci flow (see [2, Corollary 4.20] for example). Here, $\overline{\lambda}_{Ric}(x)~(x\in M)$ is the largest eigenvalue of $\mathrm{Ric}_{g}(x)$ on $T_{x}M$, and $Q_{s}(\nabla^{g}\mathrm{Ric}_{g})=O(s^{2})\cdot Q(\nabla^{g}\mathrm{Ric}_{g})$, where $Q(\nabla^{g}\mathrm{Ric}_{g})$ is a contraction of three copies of $g_{s}^{-1}$ and two copies of $\nabla^{g}\mathrm{Ric}_{g}$.
$\begin{split}&\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)^{-1}\\ &=\left(1+s\frac{R_{g}}{n}\right)^{-1}g^{-1}+s\left(1+s\frac{R_{g}}{n}\right)^{-2}g^{-1}\mathrm{Ric}_{g}g^{-1}+O(s^{2}).\end{split}$
Therefore,
$\begin{split}\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)^{-1}&\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}\left(-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)g^{-1}*\mathrm{Ric}_{g}\\ &=\frac{s^{2}\frac{R_{g}^{3}}{n^{2}}}{1+s\frac{R_{g}}{n}}+O(s^{2})\\ &=-R_{g}\left(1-s\frac{R_{g}}{n}\right)+\frac{R_{g}}{1+s\frac{R_{g}}{n}}+O(s^{2}).\end{split}$
Substituting this into the formula above, we have
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|^{2}_{1,g_{s}}\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\overline{\lambda}_{Ric}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~~~~~-R_{g}\left(\frac{s\overline{\lambda}_{Ric}}{\left(1-s\overline{\lambda}_{Ric}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~~~~~-\frac{sR_{g}}{\left(1-s\overline{\lambda}_{Ric}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\overline{\lambda}_{Ric}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2})\end{split}$ (17)
Then the desired result for the case of $s>0$ follows directly from this formula. The claim for the case of $s<0$ can be proved similarly by replacing $\overline{\lambda}_{Ric}$ with $\underline{\lambda}_{Ric}$ in the above argument. Here $\underline{\lambda}_{Ric}(x)~(x\in M)$ is the smallest eigenvalue of $\mathrm{Ric}_{g}(x)$ on $T_{x}M$
Proof of Theorem 1.4.
Let $\lambda_{Ric,1}(x)\leq\cdots\leq\lambda_{Ric,n-1}(x)\leq\lambda_{Ric,n}(x)$ be the eigenvalues of $\mathrm{Ric}_{g}(x)$ on $T_{x}M~(x\in M)$. As in the proof of Theorem 1.3, we have that for all sufficiently small $0<|s|<<1$ (so that $(1+(R_{g}/n)s)g-s\mathrm{Ric}_{g}$ is positive definite on $M$),
if $R_{g}\leq 0$ and $s>0$,
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,1}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~~~~~~~~~~-R_{g}\left(\frac{s\lambda_{Ric,1}}{\left(1-s\lambda_{Ric,1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~~~~-\frac{sR_{g}}{\left(1-s\lambda_{Ric,1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,1}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
If $R_{g}\geq 0$ and $s>0$,
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,n}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~~~~~~~~~~-R_{g}\left(\frac{s\lambda_{Ric,n}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~~~~-\frac{sR_{g}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,n}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
Then, the assertion corresponding to $s>0$ follows directly from these estimates.
For the case of $s<0$, one can prove the desired assertion by replacing $\lambda_{Ric,1},\lambda_{Ric,n}$ with $\lambda_{Ric,n},\lambda_{Ric,1}$ respectively for each case: $R_{g}\leq$ and $\geq 0$ respectively. (Of course, the second inequality sign should be the opposite for each estimate in this case.) ∎
Proof of Theorem 1.5.
From the formula (13), we have
$\begin{split}&R_{g_{s}}\cdot g_{s}\\ &=\left(R_{g}+\frac{n-1}{n}s\Delta_{g}R_{g}-s\frac{R_{g}^{2}}{n}+\frac{s}{2}\Delta_{g}R_{g}+s\|\mathrm{Ric}_{g}\|^{2}_{g}+O(s^{2})\right)\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)\\ &=R_{g}\cdot g+\left(\frac{n-1}{n}s\Delta_{g}R_{g}+\frac{s}{2}\Delta_{g}R_{g}+s\left(\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{R_{g}^{2}}{n}\right)\right)\cdot g-sR_{g}\cdot\mathrm{Ric}_{g}+O(s^{2})\end{split}$
@@
As the proof of Theorem 1.5, we have

### v4

@@ 1 Introduction
For example, the standard even-dimensional spheres satisfy the assumption of Theorem 1.1. For the case of $M=M_{0}$ and $f=\mathrm{id}_{M}$, he also gave the following type of rigidity theorem.
@@
On the other hand, the same quantity $R_{g}\cdot g$ appears in the context of Yamabe metrics.
Theorem 1.3 ([14]).
$R_{g}\cdot g\geq R_{h}\cdot h~~\mathrm{on}~M.$ (1)
Remark 1.1.
@@
By the way, a certain infinitesimal feature of Einstein metrics relates to another type of scalar curvature rigidity.
Theorem 1.4 ([17, Theorem 1.5, Remark 1.6]).
[…]
Remark 1.2.
Dahl–Kröncke [8, 17] also found a relation between stability of Einstein metrics and certain type of scalar curvature rigidity.
As noted in [10], one can apply the construction of Lohkamp [23] to see that not all metrics on $M$ are area-extremal (for this definition, see the beginning of Section 2) if $\mathrm{dim}M\geq 3$. Similarly, for each closed manifold $M^{n}$ of dimension $n\geq 3$, there is at least one Riemannian metric that is not type III scalar curvature rigid in the sense of Listing (see Definition 1.3 below). However, such an example is not given in an explicit way. That is, we can deduce the existence of such a metric but we cannot know any concrete properties involving its curvatures in general. In light of the above, this study aims to investigate relations between Listing-type of rigidity phenomena involving scalar curvature and standard metrics in various senses. In particular, we will investigate the relations between such types of scalar curvature rigidity and some infinitesimal features for constant scalar curvature metrics.
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ and a smooth function $f\in C^{\infty}(M)$ satisfying
$\begin{split}&\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}\\ &\qquad+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}-R_{g}\cdot\max_{v\in T_{x}M,\,|v|_{g}=1}\overset{\circ}{\mathrm{Ric}_{g}}\end{split}$ (2)
is positive (resp. negative) on $M$. Then, there is a small constant $s_{0}>0$ depending only on $n,M$ and $g$ such that for all $0<s\leq s_{0}$ $g_{s}=g+s\,\overset{\circ}{\mathrm{Ric}_{g}}$ (resp. $g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$) is a Riemmanian metric on $M$ and that
$R_{g_{s}}>R_{g}\cdot\|g\|^{2}_{1,g_{s}}$ (3)
on $M$. Here, $\|g\|_{1,g_{s}}:M\rightarrow[0,\infty)$ is the function on $M$ defined by
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ and a smooth function $f\in C^{\infty}(M)$ satisfying
$\begin{split}&\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}\\ &\qquad+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}-\frac{R_{g}}{2}\cdot\max_{v\in\Lambda^{2}T_{x}M\setminus\{0\}}g^{-1}\overset{\circ}{\mathrm{Ric}_{g}}(v,v),\end{split}$
where for $(x,v)=(x,v_{1}\wedge v_{2})\in\Lambda^{2}TM$,
$\begin{split}g^{-1}h(v,v)&=\mathrm{tr}\left(\begin{pmatrix}g(v_{1},v_{1})&g(v_{1},v_{2})\\ g(v_{2},v_{1})&g(v_{2},v_{2})\end{pmatrix}^{-1}\cdot\begin{pmatrix}h(v_{1},v_{1})&h(v_{1},v_{2})\\ h(v_{2},v_{1})&h(v_{2},v_{2})\end{pmatrix}\right)\\ &=\frac{1}{g(v,v)}\left(g(v_{2},v_{2})h(v_{1},v_{1})-2g(v_{1},v_{2})h(v_{1},v_{2})+g(v_{1},v_{1})h(v_{2},v_{2})\right),\end{split}$
is positive (resp. negative) on $M$. Then, there is a small constant $s_{0}>0$ depending only on $n,M$ and $g$ such that for all $0<s\leq s_{0}$ $g_{s}=g+s\,\overset{\circ}{\mathrm{Ric}_{g}}$ (resp. $g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$) is a Riemmanian metric on $M$ and that
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ and a smooth function $f\in C^{\infty}(M)$ satisfying
$\begin{split}&\left\{\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}\right.\\ &\qquad\qquad\left.+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}\right\}g+R_{g}\cdot\overset{\circ}{\mathrm{Ric}_{g}}\end{split}$
is positive (resp. negative) definite on $TM$. Then, there is a small constant $s_{0}>0$ depending only on $n,M$ and $g$ such that for all $0<s\leq s_{0}$ $g_{s}=g+s\,\overset{\circ}{\mathrm{Ric}_{g}}$ (resp. $g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$) is a Riemmanian metric on $M$ and that
$R_{g_{s}}\cdot g_{s}>R_{g}\cdot g$ (8)
on $TM$. In particular, if a metric $g$ satisfies $\mathrm{Ric}_{g}>0$ on $M$ and there is a function $u\in C^{\infty}_{+}(M)$ such that $\left\langle\mathrm{Hess}\,u,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle-\|\mathrm{Ric}_{g}\|_{g}^{2}\,u>0$ on $M$, then $g$ satisfies the above assumption.
@@
Let $(M^{n},g)$ be a closed Riemannian manifold of dimension $n\geq 2$ and a smooth function $f\in C^{\infty}(M)$ satisfying
$\begin{split}&R_{g}\left\{\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}\right.\\ &\qquad\left.+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}\right\}g+\frac{R_{g}^{2}}{2}\cdot g^{-1}\overset{\circ}{\mathrm{Ric}_{g}}\end{split}$
is positive (resp. negative) definite on $\Lambda^{2}TM$. Then, there is a small constant $s_{0}>0$ depending only on $n,M$ and $g$ such that for all $0<s\leq s_{0}$ $g_{s}=g+s\,\overset{\circ}{\mathrm{Ric}_{g}}$ (resp. $g_{s}=g-s\,\overset{\circ}{\mathrm{Ric}_{g}}$) is a Riemmanian metric on $M$ and that
$R^{2}_{g_{s}}\cdot g_{s}>R^{2}_{g}\cdot g$ (9)
@@
  • All the assumptions of the above Theorems 1.5, 1.6, 1.7 and 1.8 especially imply that $R_{g}$ is not a constant on $M$ or
@@
  • Since every Einstein metric does not satisfy any of the above assumptions (see Remark 1.4 above), our theorems above cannot be applied to Einstein metrics.
@@
This paper is organized as follows. In Section 2, we review some rigidity results for certain Riemannian metrics with positive scalar curvature and a relation between “scalar minimum functional” and an extremal metric (see Remark 2.1). In Section 3, we describe a formula that is necessary to prove our main theorems. Furthermore, we consider statements of the same type as our main theorems on compact manifolds with boundary and non-compact complete manifolds. In Section 4, we prove our main theorems. In Section 5, we give some examples that satisfy the assumptions of Theorem 1.5, 1.6, 1.7 and 1.8. In Section 6, we present some further questions related to our main theorems. In Section 7, we give a proof of the formula in Section 3.
@@ 4 Proofs of Main Theorems
[…]
@@
Let $g_{s}=g+s\,e^{f}\,\overset{\circ}{\mathrm{Ric}_{g}}$. Then, from the formula (14) and the contracted second Bianchi identity, we have
@@
$\begin{split}&R_{g_{s}}\cdot g_{s}\\ &=\left(R_{g}+se^{f}\left[\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\mathrm{Ric}_{g}\|_{g}^{2}\right.\right.\\ &\qquad\left.\left.+\frac{R^{2}_{g}}{n}+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}\right]+O(s^{2})\right)\\ &\qquad\times\left(g-s\,e^{f}\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)\\ &=R_{g}\cdot g+e^{f}\left[\left\{\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\mathrm{Ric}_{g}\|_{g}^{2}\right.\right.\\ &\qquad\left.\left.+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}\right\}g+R_{g}\cdot\mathrm{Ric}_{g}\right]+O(s^{2})\end{split}$
@@
Let $g_{s}=g+s\,e^{f}\,\overset{\circ}{\mathrm{Ric}_{g}}$ Then, as in the above proof of Theorem 1.7, we have